# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ 5 Evaluation
Table 1: Comparisons between the proposed algorithms and SOTA offline model-based RL methods on the D4RL benchmark suite. Each value represents the normalized score, as proposed in (Fu et al. (2020)), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For our algorithms, we report the average score of the final ten policy learning epochs and its standard deviation across three random seeds. Results in the last four columns are taken from the original papers (Lu et al. (2022); Yu et al. (2021); Kidambi et al. (2020); Yu et al. (2020)), respectively.
We first evaluate our algorithms on a widely-used continuous control benchmark for offline RL methods – D4RL MuJoCo (Fu et al. (2020)). The evaluation results for three types of robotic agents, each with offline datasets of four different qualities, are presented in Table 1. (1) Compared to SOTA offline MBRL methods, our algorithms achieve superior performance on nine out of twelve tasks. In terms of average performance, BA-MBRL significantly outperforms the baselines, demonstrating the effectiveness of using BAMDPs to handle model uncertainties in offline MBRL. (2) Both BA-MCTS and BA-MCTS-SL further improve upon BA-MBRL, highlighting the enhancement brought by deep search in policy learning. Notably, we apply Continuous BAMCP to only 10% of states when collecting training trajectories, while for the remaining states, we sample actions directly from the policy, i.e., $a\sim\pi(\cdot|s)$. Increasing the search ratio could further enhance policy performance at the cost of increased computation. (3) BA-MCTS-SL performs similarly to BA-MCTS, validating the effectiveness of both policy update mechanisms. However, BA-MCTS-SL struggles on Walker2d, where a warm-up training phase (using BA-MBRL) is required to establish a better initial policy. On the other hand, the advantage of the SL-based policy update is evident in the training plots of our algorithms in Figure 3, where BA-MCTS-SL exhibits much smoother learning curves compared to the other two algorithms, indicating greater robustness in model selection. (4) We further compare our algorithms with model-free offline policy learning methods, as shown in Appendix D. The performance improvement is even greater than that over model-based methods, highlighting the necessity of model-based learning. Particularly, when data quality is low, merely mimicking or staying close to the behavior policy would result in an underperforming policy. (5) We provide an ablation study in Appendix F to demonstrate the necessity of incorporating the reward penalty in offline MBRL to prevent the overexploitation of inaccurate world models. Additionally, we find that the SL-based policy update is less sensitive to model inaccuracies.
(a) hc-med-expert
(b) hc-med-replay
(c) hc-medium
(d) hc-random
(e) hp-med-expert
(f) hp-med-replay
(g) hp-medium
(h) hp-random
(i) wk-med-expert
(j) wk-med-replay
(k) wk-medium
(l) wk-random
Figure 1: Performance of Sampled EfficientZero on D4RL MuJoCo tasks. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in undiscounted episodic return as a function of the number of training samples. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively.
Figure 2: Evaluation results for the tokamak control tasks. The figure shows the change in episodic returns over training epochs for the proposed algorithms and baselines across three target tracking tasks in the nuclear fusion scenario. Solid lines represent the average performance, while shaded areas indicate the 95% confidence intervals.
For fair comparisons and real-time execution, we do not perform test-time search and adopt the same policy architecture as the baselines, i.e., a feedforward neural network, rather than an RNN that incorporates transition history as input. These alternative designs have the potential to further improve our algorithms. For implementation, we build on the codebase of Optimized (Lu et al. (2022)), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings [^9] The detailed hyperparameter setups of our algorithms are provided in Appendix C.. Therefore, we believe the performance improvements stem from the Bayesian RL framework and deep search component. Further, both components can be seamlessly integrated with other advancements in offline MBRL, such as more accurate world model learning and improved uncertainty quantification for constructing pessimistic MDPs.
MuZero also applies deep search to MBRL. To evaluate its performance on D4RL MuJoCo tasks, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al. (2021)) provided by LightZero (Niu et al. (2023)). Benchmarking results from LightZero indicate that Sampled EfficientZero, equipped with a Gaussian policy, achieves the best performance on (online) MuJoCo locomotion tasks compared to other MuZero variants. To adapt Sampled EfficientZero for offline learning, we employ the reanalyse technique proposed by (Schrittwieser et al. (2021)). The evaluation results are presented in Figure 1. For reference, the expert-level episodic returns (corresponding to scores of 100) for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively. As shown, the results are significantly worse compared to the performance of offline RL methods listed in Table 1, despite Sampled EfficientZero’s higher computational cost (detailed in Appendix E). Notably, both Sampled EfficientZero and BA-MCTS-SL rely on supervised learning for policy improvement. However, for continuous control tasks, the agent can only sample a finite number of actions at a decision point, and the search result (e.g., $\pi_{\text{ret}}$ in Algorithm 1) is a distribution over this finite set, which could be a poor approximation of the optimal action distribution. Thus, purely mimicking the search result may be less sample-efficient than policy gradient methods, as it fails to account for the continuous nature of the action space. Furthermore, world model learning is the foundation of MBRL and can be particularly challenging in continuous control and offline learning settings, where the state-action space is vast but training data is limited. Sampled EfficientZero integrates model learning and policy training into a single stage, which significantly increases the learning difficulty (compared to BA-MCTS-SL).
Finally, we evaluate our algorithms on three target tracking tasks in tokamak control. The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion, where the primary challenge lies in confining the plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions (Pironti & Walker (2005)). Tokamak control involves applying a series of direct actuators (e.g., neutral beam, ECH power, magnetic field) and indirect actuators (e.g., setting targets for the plasma shape and density) to confine the plasma to achieve a desired state or track a given target. This sophisticated physical process is an ideal test bed for our algorithms. Specifically, we use a well-trained data-driven dynamics model provided by Char et al. (2024) as a “ground truth" simulator for the nuclear fusion process during evaluation, and generate a dataset containing 725270 transitions using this model for offline RL. We select a reference shot (i.e., an episode of a fusion process) from DIII-D [^10] DIII-D is a tokamak device located in San Diego, California, operated by General Atomics., and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. These are critical quantities in tokamak control, particularly $\beta_{n}$, which serves as an economic indicator of the efficiency of nuclear fusion. The tracking tasks have a 28-dimensional state space and a 14-dimensional action space, both continuous. Moreover, these tasks are highly stochastic, as the underlying dynamics model is a probabilistic neural network and each state transition is a sample from this model. For details on the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix G. We compare our algorithms with SOTA model-free and model-based offline RL methods, specifically CQL and Optimized. The learning performance on the three tracking tasks is shown in Figure 2, where the x-axis and y-axis represent the training epochs and (negative) full-shot tracking errors, respectively. Our algorithms consistently outperform the baselines. Notably, the offline dataset does not include the reference shot or any similar, nearby shots. Therefore, restricting the policy to stay close to the behavior policy, as done in model-free methods, can be problematic. Also, learning dynamics models for MBRL is quite challenging in this nuclear fusion scenario. Our algorithms share the same ensemble of dynamics models with “Optimized" for policy learning, and the comparisons can demonstrate the superiority of Bayesian RL and deep search. Figure 2 has been smoothed for visualization [^11] The episodic return is plotted every 10 training epochs, with the y-axis representing the average value of a sliding window of length 5.. We further report the average return over the final 10 training epochs in Table 2, and the conclusions align with those from the D4RL MuJoCo tasks, showing the robustness of our proposed algorithms.
Task BA-MCTS -SL (ours) BA-MCTS (ours) BA-MBRL (ours) CQL Optimized
Temperature -21.16 $\pm$ 5.00 -23.83 $\pm$ 9.66 -29.35 $\pm$ 4.72 -59.62 $\pm$ 1.57 -83.55 $\pm$ 10.56
Rotation -14.14 $\pm$ 1.88 -19.07 $\pm$ 5.85 -31.33 $\pm$ 11.54 -85.48 $\pm$ 2.72 -71.54 $\pm$ 9.88
$\beta_{n}$37.03 $\pm$ 17.98 -18.93 $\pm$ 1.75 -23.4 $\pm$ 10.77 -36.37 $\pm$ 1.17 -57.84 $\pm$ 10.27
Average -24.11 -20.61 -28.03 -60.49 -70.98
@@
Table 2: Comparisons between the proposed algorithms and offline RL baselines on the target tracking tasks. For each algorithm, we report the average return of the final ten policy learning epochs and its standard deviation across three different random seeds.
6 Conclusion and Discussions
In this work, we propose framing offline model-based reinforcement learning (MBRL) as a Bayes Adaptive Markov Decision Process (BAMDP) to better address uncertainties in the world models learned from offline datasets. We also introduce a novel planning method for solving BAMDPs in continuous state and action spaces using Monte Carlo Tree Search. This planning process is integrated into a policy iteration framework, enabling the derivation of a policy suitable for real-time execution from the planning results. In our evaluation, we test several variants of our algorithms to separately highlight the effectiveness of Bayesian RL and deep search. Additionally, we compare two different approaches for policy updates (based on the search results) in continuous control tasks: supervised learning and policy gradient methods. Our findings demonstrate that: (1) adapting beliefs over an ensemble of world models based on experience yields more accurate model approximations for MBRL; (2) deep search improves learning performance by incorporating planning and additional computation input; and (3) while supervised-learning-based policy updates result in smoother learning curves, they may struggle in complex continuous control tasks due to their approximation of the continuous action space as a finite set of action samples. For future work, our algorithms can be further improved by integrating advancements in offline MBRL and Bayesian RL, such as Bayesian RL methods that do not rely on deep ensembles, techniques to address sparse rewards in MBRL, and more principled approaches to construct pessimistic MDPs beyond those based on ensemble discrepancy.
@@ Appendix A BAMCP
Bayes Adaptive Monte Carlo Planning (BAMCP, Guez et al. (2013)) is a sample-based online planning method, aiming to find the action $a^{*}$ that approximately maximizes the expected return at a decision point $(s,h)$ under the BAMDP. Its detailed pseudo code is shown as Algorithm 3. BAMCP has demonstrated success in solving BAMDPs with large-scale discrete state and action spaces. Its key algorithmic ideas include: (1) applying MCTS with an efficient exploration strategy – UCT (Kocsis & Szepesvári (2006)) to the BAMDP in order to simulate the outcomes of different action choices; (2) utilizing root sampling to avoid frequent Bayesian posterior updates. Specifically, the UCT rule is used for selecting actions at non-leaf nodes, i.e., $a\leftarrow\arg\max_{x}Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$, managing the tradeoff between exploration and exploitation. Root sampling refers to sampling the dynamics model only at the root node (i.e., $\theta\sim b(\cdot)$) and not adapting the belief $b(\cdot)$ according to the Bayes rule during the search process, of which the rationality is justified in the following lemma.

### v4

@@ 5.2 Benchmarking Results on D4RL MuJoCo
Table 2: Comparisons of our algorithm against SOTA offline RL methods on the D4RL benchmark. Each value represents the normalized score, as proposed in Fu et al. (2020), of the policy trained by the algorithm. For our algorithm, we report the average score of the final ten learning epochs and its standard deviation across three random seeds. To ensure a fair comparison, the results for MOBILE and CBOP are from our own runs using their publicly codebases. The results for all other baselines are taken from their respective original papers. Please check Appendix F for the references.
Data Type Environment BA-MCTS (ours) MOBILE CBOP RAMBO APE-V MAPLE Optimized
random HalfCheetah 41.45 (1.40) 39.27 (1.38) 32.95 (0.14) 40.0 29.9 41.5 31.7
random Hopper 31.42 (0.10) 7.965 (0.43) 23.79 (11.2) 21.6 31.3 10.7 12.1
random Walker2d 21.54 (0.06) 21.47 (0.03) 1.391 (0.06) 11.5 15.5 22.1 21.7
medium HalfCheetah 77.31 (1.40) 76.42 (1.00) 69.98 (0.88) 77.6 69.1 48.5 45.7
medium Hopper 103.9 (0.33) 102.5 (1.77) 98.68 (2.30) 92.8 - 44.1 69.3
medium Walker2d 88.23 (0.78) 87.12 (2.79) 93.92 (1.92) 86.9 90.3 81.3 79.7
med-replay HalfCheetah 71.24 (2.65) 64.67 (8.14) 63.36 (0.47) 68.9 64.6 69.5 58.0
med-replay Hopper 106.4 (0.53) 105.7 (1.56) 101.3 (1.74) 96.6 98.5 85.0 90.8
med-replay Walker2d 91.42 (1.54) 85.12 (10.1) 74.56 (1.29) 85.0 82.9 75.4 65.8
med-expert HalfCheetah 102.3 (2.27) 100.9 (4.00) 97.91 (5.26) 93.7 101.4 55.4 104.2
med-expert Hopper 111.8 (0.82) 111.9 (0.34) 111.2 (0.77) 83.3 105.7 95.3 105.8
med-expert Walker2d 116.0 (1.49) 115.0 (0.30) 117.6 (0.27) 68.3 110.0 107.0 97.1
Average Score 80.25 76.50 73.87 68.85 72.65 61.32 65.16
Model-Free Methods. We also compare our algorithms with a series of model-free offline policy learning methods, with detailed results presented in Appendix F.2. Our algorithms show significantly better performance, demonstrating the necessity of model-based learning in these environments, particularly when data quality is low. Notably, APE-V (see Table 2) is a model-free method built on BAMDPs, yet it shows inferior performance compared to ours in 10 out of 12 tasks. As a summary, in Table 3444MOReL, IQL, and the model-free offline RL baselines in Appendix F.2 did not report standard deviations. Additionally, IQL and DT were not evaluated on D4RL tasks with random data., we show comparisons against 19 strong model-based and model-free baselines, where BA-MCTS achieves the highest average score. (See Appendix F.3 for details of the selected baselines.) To quantify the significance of this improvement, we compute Cohen’s $d$ (Cohen, 1988) against all baselines for which standard deviations are available. The results show that the Cohen’s $d$ value is greater than 1.8 in all cases. According to Cohen’s Rule of Thumb, a value of $d>1$ indicates a very large (statistically significant) difference. This comprehensive comparison demonstrates that our proposed method establishes a new SOTA in offline RL.
Table 3: Comprehensive performance comparison on D4RL MuJoCo. Average scores for our algorithm and baselines are reported. For baselines that report standard deviations, we compute Cohen’s $d$ to quantify the statistical significance of the improvement achieved by our method. A value of $d>1$ indicates a very large and statistically significant difference.
Algorithm BA-MCTS Optimized COMBO MOReL MOPO CQL BEAR
Avg. Score 80.3 (1.1) 65.2 (5.4) 66.8 (3.4) 64.4 36.7 (11.0) 54.9 39.8
Avg. (No Random) 96.5 (1.3) - - - - - -
Cohen’s $d$ - 3.5 4.6 - 4.7 - -
Algorithm BRAC-v SAC BC MOBILE CBOP RAMBO APE-V
Avg. Score 31.4 4.1 25.1 76.5 (2.7) 73.9 (2.2) 68.9 (2.8) 72.7 (1.8)
Avg. (No Random) - - - - - - -
Cohen’s $d$ - - - 1.8 3.7 4.8 4.8
Algorithm MAPLE TD3+BC IQL SAC-N EDAC DT
Avg. Score 61.3 (1.8) 54.6 (2.4) - 76.3 (0.6) 76.0 (2.3) -
Avg. (No Random) - - 77.0 - - 74.7 (1.2)
Cohen’s $d$ 12.7 12.4 - 4.8 2.2 17.4
@@
Table 4: Comparisons on the target tracking tasks. For each algorithm, we report the average return of the final ten policy learning epochs and its standard deviation across three different random seeds.
Task BA-MCTS (Ours) BA-MCTS-SL (Ablation) BA-MBRL (Ablation) CQL Optimized
Temperature -23.83 (9.66) -21.16 (5.00) -29.35 (4.72) -59.62 (1.57) -83.55 (10.56)
Rotation -19.07 (5.85) -14.14 (1.88) -31.33 (11.54) -85.48 (2.72) -71.54 (9.88)
$\beta_{n}$ -18.93 (1.75) -37.03 (17.98) -23.40 (10.77) -36.37 (1.17) -57.84 (10.27)
Average -20.61 -24.11 -28.03 -60.49 -70.98
Search-Based Methods. MuZero also applies deep search to MBRL. To evaluate its performance on D4RL MuJoCo, we use the open-source implementation and hyperparameter configurations of Sampled EfficientZero (Ye et al., 2021) provided by LightZero (Niu et al., 2023). Benchmarking results from LightZero indicate that Sampled EfficientZero achieves the best performance on (online) MuJoCo tasks compared to other MuZero variants. To adapt Sampled EfficientZero for offline learning, we employ the reanalyse technique proposed by Schrittwieser et al. (2021). The evaluation results are presented in Figure 2 (Appendix F.4). As shown, the results are significantly worse compared to the performance of offline RL methods listed in Table 2, despite Sampled EfficientZero’s much higher computational cost. In Appendix F.4, we provide a detailed comparison of the computational costs and algorithm designs between our method and Sampled EfficientZero. World model learning is the foundation of MBRL and can be particularly challenging in continuous control and offline settings, where the state-action space is vast but training data is limited. Sampled EfficientZero integrates model learning and policy training into a single stage, which significantly increases the learning difficulty.
5.3 Ablation Study
For a principled ablation study, we reimplement BA-MCTS and its ablated variants based on the codebase of Optimized (Lu et al., 2022), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings (see Appendix D). Specifically, we first add a belief adaptation module to Optimized to obtain BA-MBRL, and then incorporate the Continuous BAMCP planner (Section 4.2) to create BA-MCTS. Finally, we replace the policy distillation module of BA-MCTS from SAC to supervised learning (as detailed in Section 4.3), yielding BA-MCTS-SL. The detailed results are presented in Table 11, and Appendices F.5 - F.6. BA-MBRL significantly outperforms Optimized, demonstrating the necessity of Bayesian model adaptation. BA-MCTS further improves upon the performance of BA-MBRL, underscoring the effectiveness of employing deep search as a powerful policy improvement operator. Finally, BA-MCTS-SL achieves performance comparable to BA-MCTS. However, BA-MCTS-SL can be less stable on certain complex tasks: it struggles on the Walker2d environments, where a warm-up phase (using BA-MBRL) is needed to establish a good initial policy for supervised learning to be effective. Finally, we conduct an ablation study to assess the necessity of the reward penalty. By removing this penalty term (i.e., setting $\lambda=0$ in Eq. (5)), we observe a consistent drop in average performance. This demonstrates that the pessimistic reward penalty is crucial for stopping the agent exploiting inaccuracies in the learned world models. Results are in Table 12 in Appendix F.7.
5.4 Application to Tokamak Control
We evaluate our algorithm on three target tracking tasks in tokamak control. The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion. We use a well-trained data-driven dynamics model provided by Char et al. (2024) as a “ground truth” simulator for the nuclear fusion process during evaluation, and generate a dataset (by this dynamics model) containing 725,270 transitions for offline RL. We select a reference shot (i.e., an episode of a fusion process) from DIII-D (a tokamak device located in San Diego), and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. The tracking tasks have a 28-dimensional state space and a 14-dimensional action space, both continuous. Moreover, these tasks are highly stochastic, as the underlying dynamics model is a probabilistic neural network and each state transition is a sample from this model. For details on the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix E.
We obtained access to the DIII-D data for a limited time window, which restricted our ability to provide evaluation results across a wide range of algorithms. For a principled comparison with the baselines, we utilized the codebase from the ablation study, where our algorithm implementations are primarily based on Optimized, to ensure that performance improvements are attributable to algorithmic design rather than implementation differences. Additionally, we included evaluation of a widely-used model-free offline RL method: CQL. The average tracking errors over the final 10 training epochs are reported in Table 4, where BA-MCTS outperforms all baselines. These results demonstrate the potential of our algorithm for application in complex, data-driven control problems.
6 Conclusion
We propose framing offline MBRL as a BAMDP to better address uncertainties in the world models learned from offline datasets. We also introduce a novel planning method for solving BAMDPs in continuous state and action spaces using MCTS. This planning process is integrated into a policy iteration framework, enabling the derivation of a policy function for real-time execution. In our evaluation, we test several variants of our algorithms to separately highlight the effectiveness of Bayesian RL and deep search. Additionally, we validate the superior performance of our approach on both standard offline RL benchmarks and challenging continuous control tasks in Nuclear Fusion.
@@ Appendix A BAMCP
Bayes Adaptive Monte Carlo Planning (BAMCP) (Guez et al., 2013) is a sample-based online planning method, aiming to find the action $a^{*}$ that approximately maximizes the expected return at a decision point $(s,h)$ under the BAMDP. Its detailed pseudo code is shown as Algorithm 3. BAMCP has demonstrated success in solving BAMDPs with large-scale discrete state and action spaces. Its key algorithmic ideas include: (1) applying MCTS with an efficient exploration strategy – UCT (Kocsis and Szepesvári, 2006) to the BAMDP in order to simulate the outcomes of different action choices; (2) utilizing root sampling to avoid frequent Bayesian posterior updates. Specifically, the UCT rule is used for selecting actions at non-leaf nodes, i.e., $a\leftarrow\arg\max_{x}Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$, managing the tradeoff between exploration and exploitation. (As shown in Algorithm 3, $N$ is the visit count, while $Q$ denotes the accumulated reward.) Root sampling refers to sampling the dynamics model only at the root node (i.e., $\theta\sim b(\cdot)$) and not adapting the belief $b(\cdot)$ according to the Bayes rule and experience during the search process, of which the rationality is justified in the following lemma.
@@
Here, $Z=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)b(\theta|h)d\theta=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)\tilde{b}(\theta|h)d\theta=1/\int_{\theta}\mathcal{P}_{\theta}(s^{\prime}|s,a)\tilde{b}(\theta|ha)d\theta$ is the normalization constant. The fifth equality in Eq. (6) holds due to the inductive hypothesis. The sixth equality is based on the fact that the choice of $a$ at each node $h$ is made independently of the sample $\theta$. As for the last equality, to experience $\theta$ at $has^{\prime}$, the sample $\theta$ needs to traverse $ha$ (with probability $\tilde{b}(\theta|ha)$) and then the state $s^{\prime}$ needs to be sampled, which is with probability $\mathcal{P}_{\theta}(s^{\prime}|s,a)$, so $\tilde{b}(\theta|has^{\prime})\propto\tilde{b}(\theta|ha)\mathcal{P}_{\theta}(s^{\prime}|s,a)$. ∎
Appendix B Consistency Proof of Continuous BAMCP
This appendix provides the detailed proof for Theorem 4.1. Our proof adapts the consistency proof methodology developed for Polynomial Upper Confidence Trees (PUCT) (Auger et al., 2013). To apply this framework, we first note that the Pessimistic Bayes-Adaptive MDP (P-BAMDP) defined in the main text can be formally treated as a standard, albeit large-state-space, MDP. Throughout this proof, we will refer to this mathematical object as the Pessimistic MDP (P-MDP). Our goal is to show that Continuous BAMCP is a consistent solver for this P-MDP.
The remainder of this appendix is organized as follows. Section B.1 formally defines the P-MDP and introduces the key notations, assumptions, and the induction property central to our proof. Section B.2 establishes the base case for our backward induction. Sections B.3 and B.4 present the two core inductive steps, showing how the consistency property propagates upward through the random and decision nodes of the search tree, respectively. Finally, Section B.5 concludes the proof.
B.1 Preamble and Definitions
P-MDP Formalization
The P-MDP that our planner solves is defined by the following components:
  • State Space $\mathcal{S}^{+}$: The augmented information state space, where each state is $(s,h)$.
  • Action Space $\mathcal{A}$: The action space of the original problem.
  • Transition Function $\mathcal{P}^{+}$: The belief-weighted transition function over information states.
  • Reward Function $\tilde{r}$: The fixed, bounded, penalized reward function from Eq. (5)
Relevant Notations
We adopt the following notations:
  • $z$: A decision node, corresponding to an information state $(s,h)$.
  • $w$: A random node, corresponding to a state-action pair $(z,a)$.
  • $V^{*}_{p}(z),V^{*}_{p}(w)$: The optimal Bellman value of node $z$ and $w$ under the P-MDP.
  • $n(z),n(w)$: The total number of visits to node $z$ and $w$ during MCTS.
  • $\hat{V}(z),\hat{V}(w)$: The empirical average return obtained at node $z$ and $w$.
  • $V(z)$: The value of a node $z$ estimated by the learned value function from the previous policy iteration.
  • $d$: The depth of a node in the search tree, integer for decision node while semi-integer for random node
Definition B.1 (Exponentially sure in $n$).
We say that some property $(P)$ depending on an integer $n$ is exponentially sure in $n$ (denoted e.s.) if there exists positive constant $C,h$ such that the probability that $(P)$ holds is at least
$1-C\exp(-hn).$
Our proof relies on the following assumptions:
Assumption B.2 (Policy Regularity).
For any decision node $z=(s,h)$ and any precision $\Delta>0$, the policy $\pi(\cdot|z)$ used to sample new actions has a non-zero probability of sampling a $\Delta$-optimal action within the current P-MDP. Formally, there exist fixed constants $\theta>0$ and $p>1$ such that,
$V^{*}_{p}(w=(z,a))\geq V^{*}_{p}(z)-\Delta\text{ with probability at least }\min(1,\theta\Delta^{p})$ (7)
Assumption B.3 (Bounded Value Function Error).
In our algorithm, when a Monte Carlo Tree Search simulation reaches a leaf node $z$ of the current search tree (i.e., a node that has not yet been expanded), we use a learned value function, $V$, to estimate its value for the backup step. This function is trained as part of the overall policy iteration framework. We assume that $V$ is an imperfect approximator of the P-MDP’s true optimal value function, $V^{*}_{p}$, and that its approximation error is uniformly bounded by a constant $\epsilon_{V}\geq 0$ over the entire state space:
$|V(z)-V^{*}_{p}(z)|\leq\epsilon_{V},\quad\forall z\in\mathcal{S}^{+}$
Furthermore, we assume the output of $V(z)$ is bounded.
The proof aims to establish the following property via backward induction:
Definition B.4 (Induction Property Cons($\gamma,f,d$)).
There exists a constant $C_{d}>0$ and a function $f_{d}(\cdot)$ such that for all decision node $z$ at depth $d$,
$|\hat{V}(z)-V^{*}_{p}(z)|\leq C_{d}n(z)^{-\gamma_{d}}+f_{d}(\epsilon_{V})\text{ e.s. in }n(z)$
and for all nodes $w$ at semi-integer depth $d+\frac{1}{2}$,
$|\hat{V}(w)-V^{*}_{p}(w)|\leq C_{d+\frac{1}{2}}n(w)^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V})\text{ e.s. in }n(w)$
where $\epsilon_{V}$ is the base approximation error of the value function (from Assumption B.3).
B.2 Base Step
The backward induction begins at the maximum search depth, $d_{max}$. We aim to establish the induction property Cons($\gamma,f,d$) for random nodes $w$ at depth $d=d_{max}-\frac{1}{2}$. As our algorithm uses a learned value function, $V(z_{i})$, to evaluate leaf nodes $z_{i}$ (children of $w$), this step relies on the value function error bound established in Assumption B.3.
Our goal is to bound the total error $|\hat{V}(w)-V^{*}_{p}(w)|$. Using the triangle inequality, we decompose this error:
$|\hat{V}(w)-V^{*}_{p}(w)|\leq|\hat{V}(w)-\mathbb{E}[V(z_{i})]|+|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$
We bound the two terms on the right-hand side separately.
(a) Bounding the Sampling Error
The first term, $|\hat{V}(w)-\mathbb{E}[V(z_{i})]|$, is the concentration error of the MCTS sampling process. Since the value function output $V(z_{i})$ is bounded (per Assumption B.3), we can apply Hoeffding’s inequality. Following the original PUCT proof (Auger et al., 2013), we set the deviation tolerance $t=n(w)^{-\frac{1}{3}}$, which gives us the following bound:
$|\hat{V}(w)-\mathbb{E}[V(z_{i})]|\leq O(n(w)^{-\frac{1}{3}})\text{ e.s. in }n(w)$
(b) Bounding the Approximation Error
The second term, $|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$, represents the inherent bias from using an approximate value function. We bound it using Assumption B.3:
$|\mathbb{E}[V(z_{i})]-V^{*}_{p}(w)|$ $=|\mathbb{E}[V(z_{i})]-\mathbb{E}[V^{*}_{p}(z_{i})]|$
$\leq\mathbb{E}[|V(z_{i})-V^{*}_{p}(z_{i})|]$
$\leq\mathbb{E}[\epsilon_{V}]=\epsilon_{V}$
Conclusion for the Base Step
By combining the bounds for the sampling error and the approximation error, we obtain the bound on the total error:
$|\hat{V}(w)-V^{*}_{p}(w)|\leq O(n(w)^{-\frac{1}{3}})+\epsilon_{V}\text{ e.s. in }n(w)$
This result precisely matches the form required by our revised induction property, Cons($\gamma,f,d$). We have successfully shown that the property holds for the base case at depth $d=d_{max}-\frac{1}{2}$, with parameters:
  • $\gamma_{d_{max}-\frac{1}{2}}=\frac{1}{3}$
  • $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$
B.3 Inductive Step: From Decision to Random Nodes
In this section, we perform the first inductive step, moving from decision nodes up to random nodes. We assume the induction property Cons($\gamma,f,d$) holds for all decision nodes $z$ at depth $d$. Our objective is to prove that, as a consequence, the property also holds for their parent nodes, which are the random nodes $w$ at depth $d-\frac{1}{2}$.
A key distinction in our algorithm is the special handling of the very first action sampled from a decision node. This requires us to categorize all random nodes $w$ (at depth $d-\frac{1}{2}$) based on their history:
  1. 1.
    Category 1: First-Child Nodes ($w_{0}$): A random node that was the first child ever created from its parent decision node. Its initial value estimate was derived directly from our learned value function, $V$.
  2. 2.
    Category 2: Standard Nodes ($w_{j},j>0$): All other random nodes. Their value estimates have always been computed through the standard MCTS backup process.
We will analyze these two categories separately to show that the induction property holds universally.
B.3.1 Analysis of Standard Random Nodes ($w_{j},j>0$)
For a standard random node $w$ in this category, the proof structure closely follows Section 5 of Auger et al. (2013). We aim to bound the total error $|\hat{V}(w)-V^{*}_{p}(w)|$.
Let $i_{0}=\lfloor n(w)^{\alpha}\rfloor$ be the newest child of $w$ (a decision node at depth $d$), and let $r=i_{0}-1$ be the number of “mature” children.(i.e., all children except for the newest). Let $R_{w}(i)=r_{i}+\gamma\hat{V}(i)$ denote the return through child $i$. We decompose the total error into four terms: